Speed of a motor launch with respect to still water is 70m/s.Speed of stream is 3m/s.When the launch begins to travel upstream a float was dropped from it.The launch travel 4.2km upstream,turned about and caughtup with the float.How long is it before the launch reaches the floor.??

Grade:

1 Answers

Laveena Wadhwani
2 years ago
Relative speed of motor launch w.r.t still water is 7m/s,but speed of still water is 0
So,velocity of motar boat is 7m/s
Now during upstream speed of boat is opposite of stream ,so speed during upstream is 7-3=4 m/ s
And during downstream both work in same direction so,7+3=10 m/s
Now the float dropped ,
Now, during upstream distance covered is 4.2 km= 4200m
 So time is = 4200/4  =1050s
Distance covered by float = speed × time = speed that of stream × 1050= 3 ×1050=3150
Total distance traveled is 3150+4200= 7350m
Now ,real speed is 7m/ s,
So t= 7350/ 7= 1050s,
Total time = 1050 +1050= 2100s= 35 min